PART FIVE Case Problem 1. Byron Bass is a commercial real estate broker who also has a keen eye for personal investment opportunities. He has an opportunity to buy Academic Arms, an apartment building that caters to students. He estimates that he can acquire the asset for $1.3 million with purchase costs of 2 percent and that the seller will take back a $1.2 million first mortgage note with the annual debt service and principal payments reflected on the following schedule of projected cash flows. Bass plans to sell the project when the loan balancıe falls due at the end of the seventh year Bass has researched Academic Arms and has generated the following sched- ules, in which he has considerable confidence. Academic Arms Apartments Projected After-Tax Cash Flows from Operations Year 1 Year 2 Year 3 Year 4 Year 5 Year 6Year 7 $286.442 $295,035 $303,877 $313,004 $322.393 $332,067 $342.029 $128,899 $132.766 $136,739 $140,852 $145,077 $149,430 $153.913 Effective gross income Less: Operating expenses "15Z543 6220g 167.198 172 152 .177316 」8263Z .208.116 Net operating income Less: Debt service Before-tax cash flow Less: Depreciation Plus: Principal paid Taxable income (loss)" 126370 126370 12637Ω 126.370 126370 lanazo 126370 $2,529 S 6,396 $ 10.369 $ 14,482 $18,707 $23.060 $27.543 36,967 38,57538,575 38,575 38,575 38,575 36.967 66717.3698.1418.993 9.935 10.975 12.124 $(27,767) $(24,810) $(20,065) (15,100) (9,933) $ (4,540) 2,700 Times: Tax rate Tax (tax saving) Before-tax cash flow 28282828 28 28 28 2,5296,39610,369 14,482 18,707 23,060 27,543 17.775) 6.947) 15.618) 14.220 L2rn 11271) -156 Less: Tax (Pius. Savings) After-tax cash flow 'Bass has adequate passive income to offset these losses. 281
Tax Consequences of hip 282 PART FIVE Measures of Investment Academic Ams Apartments Antcipaofo Disposal (at the end of seventh year ted Income of o $1,700,000 s 504,809 Sales price ss: Adjusted tax basis before sale Taxable gain Income tax on sale: Tax on recapture of straight Toa er25x$266.809) Taxon long-term capitalgain (.15 × S238. Total tax on sale sh Flow from Disposa Academic Arms Apartments Anticipated (at the end of seventh year of ownership $1,700,000 Sales price Less: Income taxes s102,402 136,000 ransaction costs (.08 x $1,700,000) Mortgage balance 1135.7921374.194 S 325.806 After-tax cash flow Required: a. Using the first-year operating forecast, compute: Gross income multiplier (using effective gross income) Net income multiplier Operating ratio Break even, or default, ratio Debt coverage ratio Overall capitalization rate Equity dividend rate Cash-on-cash return rate, discount the expected after-tax cash flows from b. Using a 12 percent the this investment and determine thefl pecteda Present value Net present value Profitability index Investment value ernal rate of return. e. Determine the project's anticipateg
283 Case Problem 2. Please refer to the Case Problems for Part Four Assignments 1 and 2 of that case required you to develop projections of after-tax cash flows from opera- tions and from disposal of the St. George Apartments. If you have not com- pleted that assignment, please do so now. Using your after-tax projections, compute the following: a. The present value of the opportunity, discounting at 14 percent b. The investment value, discounting at 14 percent c. The internal rate of return
Answer #1

a)

1. Gross Income Multiplier

= Selling value of the project / Effective Gross income

= 1200,000 / 286,442 = 4.19

2. Net Income Multiplier

= Selling value of the project / Net Operating income

= 1200,000 / 128,899 = 9.31

3. Operating Ratio

= Operating Expense / Effective Gross Income * 100

= 157,543 / 286,442 * 100 = 55%

4. Break-even or Default Ratio

= Total expenses / Effective Gross Income * 100

Where Total expenses =

Operating Expenses 157,543
Debt Service 126,370
Depreciation 36,967
Total Expenses without taxes 320,880

Alternatively, total expenses can be calculated by including tax outflow

=320,880 / 286,442 * 100 = 112.02 %

5. Debt - coverage ratio

= Net Operating Income / Debt service

= 128,899 / 126,370 = 1.02

6. Over-all Capitalization Rate

= Net Operating Income / Value of Investment

= 128,899 / 1326,000 * 100 = 9.72 %

Cost        13,00,000
Purchase cost @ 2%             26,000
Total Cost of Investment        13,26,000

7. Equity dividend rate

= Before tax cash flows / Equity Investment

=2529 / 1326,000 * 100 = 0.19 %

8. Cash on cash Return rate

= cash flow (the net operating income) (before tax) / cash initially invested * 100

= 128,899 / 1326,000 * 100 = 9.72 %

b)

Year After tax cash-flow Discounting rate @ 12%
1 10304 0.89 9200
2 13343 0.80 10637
3 15987 0.71 11379
4 18710 0.64 11891
5 21488 0.57 12193
6 24331 0.51 12327
7 26787 0.45 12117
7 1597598 0.45 722672
Total Present value 802416
Cost of the project 1326000
Net present value -523584
Profitability index = Total Present value / Cost of the project 0.61
Investment value = Present value of future cash flow 802416
Cash flow of year 7 is Cash received from the project - Tax expenses =1700000-102402

c)

Rate of return 1.04
Year After tax cash-flow Discounting rate @ IRI
1 10304 0.96 9910
2 13343 0.93 12343
3 15987 0.89 14223
4 18710 0.86 16009
5 21488 0.82 17684
6 24331 0.79 19258
7 26787 0.76 20392
7 1597598 0.76 1216181
Total Present value 1326000
Cost of the project 1326000
Net present value 0

Internal Rate of return will be 4% where NPV becomes Zero

Assignment 2

Rate of return 1.14
Year After tax cash-flow Discounting rate @ 14%
1 10304 0.88 9039
2 13343 0.77 10267
3 15987 0.67 10791
4 18710 0.59 11078
5 21488 0.52 11160
6 24331 0.46 11085
7 26787 0.40 10705
7 1597598 0.40 638460
Total Present value 712584
Cost of the project 1326000
Net present value -613416

Value of investment = 712,584

Internal rate of return will be same as option (c) i.e., 4%